OpenStudy (anonymous):

The data from 200 machine parts are summarized as follows: Depth of Bore Edge Condition above target below target Coarse 15 10 Moderate 25 20 smooth 50 80 a) What is the probability that a part selected has a moderate edge condition and a below-target bore depth? b) What is the probability that a part selected has a moderate edge condition or a below target bore depth? c) What is the probability that a part selected does not have a moderate edge condition

OpenStudy (anonymous):

Let's denote some events: \(C\): coarse edge \(M\): moderate edge \(S\): smooth edge \(A\): above target \(B\): below target a) Find \(P(M\cap B)\). This is given by the ratio of the number of machine parts that have both qualities (25) to the total (200), or \(\dfrac{25}{200}=\dfrac{1}{8}\). b) Find \(P(M\cup B)\). The probability of a union of events is given by the following formula: \[P(M\cup B)=P(M)+P(B)-P(M\cap B)\] As before, each probability is a ratio of the parts that satisfy a given characteristic to the total number of parts. c) Find \(P(\overline{M})\), or \(P(M^C)\), where \(\overline{M}\) and \(M^C\) denote the same thing: the complement of \(M\), i.e. the event that \(M\) does not occur. \[P(\overline{M})=1-P(M)\]

OpenStudy (anonymous):

in part a) the intersection of M with B should be 20 not 25 am i right @SithsAndGiggles

OpenStudy (anonymous):

b) -- P(M)=45/200 P(B)=110/200 and P(MUB)=27/40

OpenStudy (anonymous):

@yama_aryayee, correct, I misread my own table :)